OpenStudy (anonymous):

5) What is the volume of a ball if the diameter if the ball is 6 inches? Leave your answer in terms of π. A) V = 12 π cubic inches B) V = 36 π cubic inches C) V = 288 π cubic inches D) V = 27.75 π cubic inches

OpenStudy (anonymous):

Volume of spheres is 4/3 pi r cubed.

OpenStudy (anonymous):

I got B but I don't know if its right.

OpenStudy (anonymous):

V = 4/3*pi*3^3 = 4/3*pi*27 = 36*pi B is the correct answer.(: